y is directly proportional to x^2. When y = 5, x = 10.
Write down the equation of proportionality with full working

Found 2 solutions by Fombitz, Theo:
Answer by Fombitz(32388)
Answer by Theo(13342)

the direct proportion formula is y = kx.

when x = x^2, the formula becomes y = kx^2.

when y = 5, x = 10 is the statement.

you can use this to solve for k.

you get y = kx^2 becomes:
5 = k*10^2 which becomes:
5 = k*100

solve for k to get k = 5/100 = 1/20 = .05

now that you know the value of k, you can use that value to solve all other problems involving this proportion.

for example:

when y = 50, what is the value of x^2?

you know the value of k is equal to .05.

the formula of y = kx^2 becomes y = .05*x^2

y = 50 so the formula becomes 50 = .05*x^2

divide both sides of this equation by .05 to get 50/.05 = x^2

simplify to get 1000 = x^2

solve for x to get x = sqrt(1000)

the equation becomes 50 = .05 * (sqrt(1000))^2 which becomes 50 = .05 * 1000 which becomes 50 = 50 which is true.

here'e another example:

what's the value of y when x = 25?

the formula is y = kx^2.

k is equal to .05.

when x = 25, the formula becomes y = .05 * 25^2.

simplify to get y = .05 * 625

simplify further to get y = 31.25

the equation becomes 31.25 = .05 * 25^2 which becomes 31.25 = .05*625 which becomes 31.25 = 31.25 which is true.

since these are direct proportions, a/b^2 should be equal to c/d^2 which should be equal to e/f^2

a = 5
b^2 = 100

c = 50
d^2 = 1000

e = 31.25
f^2 = 625

a/b^2 should be equal to c/d^2
replace a/b^2 and c/d^2 with their respective values and you get 5/100 should be equal to 50/1000
cross multiply to get 5*1000 = 100*50
simplify to get 5000 = 5000
this is true.

c/d^2 should be equal to e/f^2
replace c/d^2 and e/f^2 with their respective values andyou get 50/1000 should be equal to 31.25/625
cross multiply to get 50*625 = 1000*31.25
simplify to get 31250 = 31250
this is true.

based on the law of equations that states if x = y and y = z, then x = z, you also get a/b^2 = e/f^2, which will also be true.

that's how the value of k is used to solve direct proportion type problems.

y = kx.

when x = x^2, the formula becomes y = kx^2.

when you know the value of y and and x, you can solve for the value of x^2 and then solve for the value of k.

once you know the value of k, you can solve other problems using the same formula of y = kx^2.

if you know the value of y and k, you can solve for the value of x^2 and then solve for the value of x.

if you know the value of of x and k, you can solve for the value of x^2 and then solve for the value of y